Southwest Sands currently has 24,000 shares of stock outstanding. It is considering issuing $140,000 of debt at an interest rate of 7.7 percent. The break-even level of EBIT between these two capital structure options is $78,000. How many shares of stock will be repurchased if the company undergoes the recapitalization? Ignore taxes.

Multiple choice...

2,606.15 shares

3,593.33 shares

2,843.08 shares

3,316.92 shares

3,151.08 shares
Answer #1

At the break-even EBIT level:

[EBIT/Original No. of Shares] = [(EBIT - Interest) / New No. of Shares]

[$78,000 / 24,000] = [{$78,000 - ($140,000 x 0.077)} / X]

$3.25 = $67,220 / X

X = $67,220 / $3.25 = 20,683.08

Shares Repurchased = Original No. of Shares - New No. of Shares

= 24,000 - 20,683.08 = 3,316.92 shares

Hence, Option "D" is correct.